A SPECTROSCOPIC STUDY OF THE LINEAR-BENT ELECTRONIC TRANSITIONS OF JET-COOLED BCl_{2} AND HBCl

\underline{RAMYA NAGARAJAN}, JIE YANG and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Laser induced fluorescence (LIF) and single vibronic level emission spectra (SVL) of jet-cooled BCl_{2} and HBCl have been measured. The radicals were produced in a pulsed electric discharge of a mixture of BCl_{3}/Ar and BCl_{3}/H_{2}/Ar, respectively. The LIF spectra of both radicals are congested due to overlapping bands from the boron and chlorine isotopes. In addition, the ground and first excited states are the two Renner-Teller components of a ^{2}Π state split by a strong vibronic interaction. The Franck-Condon profile of the Ã^{2}B_{1} - X̃^{2}A_{1} band system of BCl_{2} and the Ã^{2}A′′ - X̃^{2}A′ system of HBCl are characteristic of linear-bent excitations. Excited state bending progressions have been identified in both species using the LIF-sync scan procedure in which the monochromator is offset by a value corresponding to a ground state fundamental frequency of the target molecule and scanned simultaneously with the dye laser. LIF spectra of individual isotopes can thus be recorded. Ground state vibrational frequencies have been deduced from SVL spectra. The emission spectra for BCl_{2}, are dominated by progressions in the symmetric stretching (ν_{1}) and bending (ν_{2}) modes. In the case of HBCl, progressions in the bending (ν_{2}) and BCl stretching (ν_{3}) modes were observed.